# One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks

### v1

@@ Abstract
Language is not monolithic. While many benchmarks are used as proxies to systematically estimate Large Language Models’ (LLM) performance in real-life tasks, they tend to ignore the nuances of within-language variation and thus fail to model the experience of speakers of minority dialects. Focusing on African American Vernacular English (AAVE), we present the first study on LLMs’ fairness and robustness to a dialect in canonical reasoning tasks (algorithm, math, logic, and comprehensive reasoning). We hire AAVE speakers, including experts with computer science backgrounds, to rewrite seven popular benchmarks, such as HumanEval and GSM8K. The result of this effort is ReDial, a dialectal benchmark comprising $1.2K+$ parallel query pairs in Standardized English and AAVE. We use ReDial to evaluate state-of-the-art LLMs, including GPT-4o/4/3.5-turbo, LLaMA-3.1/3, Mistral, and Phi-3. We find that, compared to Standardized English, almost all of these widely used models show significant brittleness and unfairness to queries in AAVE. Furthermore, AAVE queries can degrade performance more substantially than misspelled texts in Standardized English, even when LLMs are more familiar with the AAVE queries. Finally, asking models to rephrase questions in Standardized English does not close the performance gap but generally introduces higher costs. Overall, our findings indicate that LLMs provide unfair service to dialect users in complex reasoning tasks. Code can be found at https://github.com/fangru-lin/redial_dialect_robustness_fairness.git.
@@ 1 Introduction
Over the last few decades, linguistics has firmly established that language varies along different external dimensions such as geography, age, and gender, dialectal variation being among the most perspicuous manifestations (Chambers & Trudgill, 1998). Speakers of ‘non-standard’ dialects are known to experience implicit and explicit forms of discrimination in everyday situations, including housing, education, work, and criminal justice (Baugh, 2005; Adger et al., 2014; Rickford & King, 2016; Drożdżowicz & Peled, 2024). As Large Language Models (LLMs) are increasingly employed as a service and by a rapidly growing user base (Milmo, 2023; La Malfa et al., 2024), it is vital to understand the service quality that they provide to different groups and demographics.
@@
In this work, we examine LLMs’ dialect robustness and fairness. Previous studies have shown that language models exhibit biases to dialect prompts in tasks such as hate speech detection and reading comprehension (Sap et al., 2019; Ziems et al., 2023), as well as making judgments about employability and criminal justice (Hofmann et al., 2024). Equally relevant, yet less studied, are tasks that require reasoning abilities for problem-solving, decision-making, and critical thinking (Wason, 1972; Huth, 2004; Huang & Chang, 2022; Qiao et al., 2022). For instance, algorithm-related tasks (e.g., generation, debugging, etc.) figure prominently in real user queries, as reflected by their first place on the ArenaHard quality board (Li et al., 2024) and their third place on the WildChat frequency board (Zhao et al., 2024). However, existing dialectal benchmarks (e.g., Ziems et al., 2023) do not cover these tasks, and current popular reasoning benchmarks such as HumanEval (Chen et al., 2021) and GSM8K (Cobbe et al., 2021) are constructed in Standardized English. It is thus unclear whether LLMs are fair when responding to reasoning tasks expressed in ‘non-standard’ English dialects. Moreover, dialect queries can also be used to test LLMs’ robustness. Adversarial robustness provides a consolidated framework to test LLMs on slight variations of existing tasks (Moradi & Samwald, 2021; Jin et al., 2023). In this sense, dialects reformulate a problem while maintaining its semantics, i.e., they test what has been referred to as semantic robustness (Malfa & Kwiatkowska, 2022).
In this work, we present the first study on evaluating LLMs in reasoning tasks expressed in African American Vernacular English (AAVE), with the objective to evaluate LLMs’ fairness and robustness towards a dialect. We choose AAVE since around $33$ million people worldwide and approximately $80\%$ of African Americans in the United States speak AAVE, with reports of discriminative behaviors in various scenarios (Lippi-Green, 1997; Purnell et al., 1999; Massey & Lundy, 2001; Grogger, 2011; Rickford & King, 2016). Our study aims to understand whether LLMs hold biases against AAVE speakers in reasoning tasks. Previous approaches in creating AAVE benchmarks from existing Standardized English data either (i) primarily use validated lexical and morphosyntactic transformation rules (Ziems et al., 2022; 2023), which fail to capture highly context-dependent nuances of dialects, or (ii) rely on LLMs as translators (Gupta et al., 2024), which may have the very biases that our research wants to unveil (Fleisig et al., 2024; Smith et al., 2024). Therefore, we hire human AAVE speakers to rewrite instances of seven popular benchmarks to AAVE, including HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and GSM8K (Cobbe et al., 2021) (see Section 2.1 for the complete list of datasets).
@@
Figure 1: ReDial is a dialect reasoning benchmark composed of $1.2K+$ Standardized English-AAVE parallel queries. Its source data comes from existing benchmarks in Standardized English. AAVE speakers are hired to rewrite each instance in their dialect but preserve their original intent, meaning, and ground truth output label to form their AAVE counterparts.
We build and release the first end-to-end human-written Standardized English-AAVE parallel benchmark called ReDial (Section 2, examples in Figure 1 and Appendix A.2). ReDial contains more than $1.2K$ Standardized English-AAVE prompt pairs, covering four fundamental reasoning tasks, namely algorithm, math, logic, and comprehensive reasoning (i.e., tasks requiring the composition of the other three reasoning skills). To the best of our knowledge, our dataset is the first high-quality reasoning dataset with parallel prompts of Standardized English and a dialect annotated end-to-end by dialect speakers. Unlike benchmarks using LLMs as judges, which are subjective to their internal biases (Zheng et al., 2023; Chen et al., 2024; Shi et al., 2024), ReDial offers an objective measure as judged by ground truth labels. It enables an easy, objective, and scalable way to report on the dialect fairness and robustness of LLMs as we keep the labels and evaluation process unchanged from the standard pipelines. We consider this dataset an important step toward revealing the robustness and fairness of state-of-the-art (SotA) LLMs for dialect users.
We use ReDial to benchmark GPT-4o, GPT-4, LLaMA-3.1-70B-Instruct, and several other widely used SotA LLMs (Section 3). We discover that almost all LLMs suffer from significant performance drops for AAVE instances, despite the fact that they are semantically equivalent to their standardized counterparts. All models except GPT-4o and LLaMA-3.1-70B-Instruct have a pass rate of less than or similar to $0.6$ in AAVE, even with Chain-of-thought prompting (CoT; Kojima et al., 2022; Wei et al., 2022), while the best pass rate in Standardized English is 0.832.
We further conduct an extensive analysis of the potential reasons for this performance gap (Section 4). We show that the skewness of dialect training data does not explain the whole picture, as large-scale LLMs have more difficulties in AAVE than misspelled Standardized English prompts, the latter of which LLMs are even less familiar with in the measurement of perplexity. This indicates that naively acquainting LLMs with AAVE by data augmentation might not be helpful for dialect robustness and fairness. Further, the performance gap cannot be easily closed by simple standardization: prompting LLMs to paraphrase AAVE in a standardized introduces higher costs, but cannot reach the Standardized English prompt performance. These findings point to the conclusion that LLMs are unfair and brittle to dialects and that the problem cannot be easily mitigated.
To summarize, the main contributions of this work are as follows:
  1. 1.
    We release ReDial, the first high-quality, human-annotated AAVE-Standardized English parallel dataset in four canonical reasoning tasks, comprising seven popular benchmarks.
  2. 2.
    We evaluate several SotA LLMs and show that they are significantly more brittle and unfair to AAVE prompts than their Standardized English counterparts, even with CoT.
  3. 3.
    Compared to misspelled Standardized English prompts of even higher perplexities, large-scale LLMs are more brittle to AAVE, which means that naive data augmentation might not solve the problem. We further find that prompting LLMs to rephrase a problem in Standardized English does not close the gap, either, but tends to introduce higher costs.
The the paper is organized as follows. We introduce ReDial in Section 2, and describe the benchmarking experiment and corresponding results in Section 3. We conduct extensive analysis in Section 4 and review related work in Section 5. Finally, we conclude the paper and discuss limitations, ethic statement, and reproducibility statement in Sections 6 to 9.
2 Dataset
Category Algorithm (19.7%) Logic (29.8%) Math (25.8%) Comprehensive (24.7%) Total
@@ 2 Dataset
Table 1: ReDial contains tasks for four categories, drawn from seven data sources. Percentage points in brackets for the categories indicate the proportion of corresponding data points in ReDial. In total, ReDial consists of $1,216$ fully-annotated parallel prompts.
In this section, we introduce ReDial (Reasoning with Dialect Queries), a benchmark of more than $1.2K$ parallel Standardized English-AAVE query pairs (see a distribution overview in Table 1 and examples in Figure 1 and Appendix A.2). Following Zhu et al. (2023a), ReDial benchmarks four canonical reasoning tasks, namely algorithm, logic, math, and comprehensive reasoning. The task formulation is linguistically diverse, addresses cornerstone problems in human reasoning, and is of particular interest as it is challenging for LLMs.
In Section 2.1, we present more details about source data collection. In Section 2.2, we describe the annotation and validation process that we used to ensure that the data is of high quality.
@@ 2.1 Data Sourcing
To obtain a highly curated dataset, we sample from seven widely used and established benchmarks. For each dataset, we report the key references, a description of the task, and the sample data size. We further provide example instances in Appendix A.1.
Algorithm HumanEval (Chen et al., 2021) contains 164 human-written instances as code completion tasks. We adopt the paradigm from InstructHumanEval [^1] https://huggingface.co/datasets/codeparrot/instructhumaneval to convert code completion headings to instruction-following style natural language queries and include all of them in our benchmark.
Algorithm MBPP Austin et al. (2021) contains $1000$ code generation queries. We include $150$ randomly sampled data points from its sanitized test instances (Liu et al., 2023).
Math GSM8K (Cobbe et al., 2021) is a dataset of graduate school math questions written in natural language. It contains $8.79K$ instances in total. We randomly sample $150$ instances from its test set.
Math SVAMP (Patel et al., 2021) contains $1000$ instances of elementary-school math problems written in natural language. We randomly sample $150$ instances from its test set.
Logic LogicBench (Parmar et al., 2024) is a benchmark of logic questions written in natural language. It contains logic questions of multi-choice and binary classification formats. We sample $100$ instances from binary and multi-choice questions each, resulting in 200 instances in total.
Logic Folio (original+perturbed) (Han et al., 2022; Wu et al., 2023) Original Folio is a manually curated logic benchmark written by students in computer science in natural language. We select $81$ instances with their manually perturbed versions from Wu et al. (2023), resulting in 162 instances in total.
Comprehensive AsyncHow (Lin et al., 2024) is a comprehensive reasoning benchmark in efficient planning with constraints. LLMs need to derive a dependency graph given natural language description (i.e., logic), find different possible paths in the graph (i.e., algorithm), and then calculate and compare the time needed for these paths (i.e., math) to reach the correct answer. We use this dataset to study whether LLMs’ robustness is dependent on compositionality. We conduct stratified sampling according to the dataset’s complexity metric and obtain $240$ instances in total.
With data points from these sources, we construct a systematic reasoning benchmark with curated data. Then, we hire AAVE speakers to rewrite these data points in their dialect.
2.2 AAVE Annotation and Quality Check
@@
Figure 2: Annotation and cross-validation of ReDial instances. We first sample instances from datasets of four canonical reasoning tasks to compose the source data, then we hire AAVE speakers to rewrite the instances in their dialect. To ensure the high quality of the rewritten data, we conduct naturalness check by AAVE speakers and correctness check by non-AAVE speakers and LLMs. We reannotate instances that do not pass the quality checks and iterate the process until the data meet our criteria. Finally, we combine the source data and AAVE rewriting to obtain a high-quality parallel reasoning dataset ReDial.
We conduct a careful data annotation and quality check pipeline, which we schematize in Figure 2 and detail below.
Annotation. We hire AAVE speakers and instruct them to rewrite each instance by making them sound natural to them, but also preserve the essential information so that ground truth labels stay unchanged (e.g., it is allowed to turn “two” into $2$, and vice-versa, but not to alter/delete numerical quantities). For algorithm tasks that require an understanding of code to keep the semantics, we specifically hire expert AAVE annotators with computer science backgrounds. [^2] Please refer to Appendix A.3 for annotators’ compensation, qualification, and other guideline details.
@@
Validation. To ensure the quality of the annotation, we conduct careful validations to ensure its naturalness and correctness. First, to ensure naturalness, we ask annotators to cross-check and edit each others’ annotations to make sure that the annotations are natural to AAVE speakers. Second, to ensure correctness, we conduct both manual and automatic checks by non-AAVE speakers and LLMs. We first have non-AAVE speakers manually check whether the rewriting maintains the essential information and send the invalid instances back to AAVE speakers for reannotation. We conduct a sanity check with GPT-4o for the correctness of rewriting (details in Appendix A.4). We manually check data that GPT-4o flags as invalid to see if all essential information is preserved: we stress that in this round no instance is rejected solely based on the LLM’s judgment. We return invalid instances to AAVE speakers for correction and iterate the process until all the data passes the check.
After this process, we obtain a high-quality, human-annotated dataset ReDial with more than $1.2K$ Standard English-AAVE parallel prompts in four canonical reasoning tasks. ReDial is the first benchmark of its kind and enables easy testing and analysis of LLMs’ dialect fairness and robustness in reasoning tasks. In the rest of the paper, we will refer to the Standardized English part of ReDial as Standardized ReDial, and its AAVE part as AAVE ReDial.
3 Experiment
In this section, we benchmark several SotA LLMs on the parallel prompts from ReDial. We report experiment setting in Section 3.1 and results in Section 3.2.

### v2

@@ Abstract
Language is not monolithic. While benchmarks, including those designed for multiple languages, are often used as proxies to evaluate the performance of Large Language Models (LLMs), they tend to overlook the nuances of within-language variation, and thus fail to model the experience of speakers of non-standard dialects. Focusing on African American Vernacular English (AAVE), we present the first study aimed at objectively assessing the fairness and robustness of LLMs in handling dialects in canonical reasoning tasks, including algorithm, math, logic, and integrated reasoning. We introduce ReDial (Reasoning with Dialect Queries), a benchmark containing 1.2K+ parallel query pairs in Standardized English and AAVE. We hire AAVE speakers, including experts with computer science backgrounds, to rewrite seven popular benchmarks, such as HumanEval and GSM8K. With ReDial, we evaluate widely used LLMs, including GPT, Claude, Llama, Mistral, and the Phi model families. Our findings reveal that almost all of these widely used models show significant brittleness and unfairness to queries in AAVE. Our work establishes a systematic and objective framework for analyzing LLM bias in dialectal queries. Moreover, it highlights how mainstream LLMs provide unfair service to dialect speakers in reasoning tasks, laying a critical foundation for relevant future research. [^1] Code and data can be accessed here.
One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks
Fangru Lin1thanks: Work done during internship at Microsoft Research. Contact: fangru.lin@ling-phil.ox.ac.uk, shaoguang.mao@microsoft.com, Shaoguang Mao2, Emanuele La Malfa1,3, Valentin Hofmann4,5 Adrian de Wynter6,7, Xun Wang6, Si-Qing Chen6 Michael Wooldridge1,3, Janet B. Pierrehumbert1, Furu Wei2 1University of Oxford, 2Microsoft Research, 3Alan Turing Institute, 4Allen Institute for AI, 5University of Washington, 6Microsoft Corporation, 7University of York
@@ 1 Introduction
Over the last few decades, linguistic research has firmly established that language naturally varies along social, geographic, and demographic dimensions (Chambers and Trudgill, 1998). Such dialectal variation is one of the most salient forms of linguistic diversity. Speakers of “non-standard” dialects are known to experience implicit and explicit discrimination in everyday situations, including housing, education, employment, and the criminal justice system (Baugh, 2005; Adger et al., 2014; Rickford and King, 2016; Drożdżowicz and Peled, 2024). As Large Language Models (LLMs) increasingly serve a broad and rapidly expanding user base (Milmo, 2023; La Malfa et al., 2024), it is critical to understand how they interact with diverse linguistic communities.
@@
In this work, we examine LLMs’ dialect robustness and fairness. For robustness, adversarial robustness provides a consolidated framework to test LLMs on slight variations of existing tasks (Moradi and Samwald, 2021; Jin et al., 2023). Dialects reformulate a problem while maintaining its semantics, i.e., they test what has been referred to as semantic robustness (Malfa and Kwiatkowska, 2022). For fairness, recent research has demonstrated that LLMs exhibit biases against non-standard dialect queries, predominantly assessed in language and social analysis tasks (Sap et al., 2019; Ziems et al., 2023; Hofmann et al., 2024). Equally relevant, yet less studied, are tasks that require reasoning abilities for problem-solving, decision-making, and critical thinking (Wason, 1972; Huth, 2004; Huang and Chang, 2022; Qiao et al., 2022). For instance, algorithm-related tasks (e.g., generation, debugging, etc.) figure prominently in real user queries, as reflected by their first place on the ArenaHard quality board (Li et al., 2024) and their third place on the WildChat frequency board (Zhao et al., 2024).
@@
Figure 1: ReDial is a dialect reasoning benchmark composed of $1,200$+ SE-AAVE parallel queries. Its source data comes from existing benchmarks in SE. AAVE speakers are hired to rewrite each instance in their dialect but preserve their original intent, meaning, and ground truth output label to form their AAVE counterparts.
However, existing dialectal benchmarks (e.g., Ziems et al., 2023) do not cover these tasks, and current popular reasoning benchmarks such as HumanEval (Chen et al., 2021) and GSM8K (Cobbe et al., 2021) are constructed in Standardized English (SE). This could disadvantage dialect speakers in real-world applications like educational assessment (González-Calatayud et al., 2021), personalized recommendation (Kantharuban et al., 2024), and even multimodal tasks (e.g., voice assistant) Martin and Wright (2022), ultimately forcing them to shift their language styles to SE (Cunningham et al., 2024) in order to access the full benefits of modern technologies.
We present the first study that systematically and objectively evaluates LLM fairness and robustness in reasoning tasks expressed in a non-standard dialect. We choose AAVE since around $33$ million people worldwide and approximately $80\%$ of African Americans in the United States speak AAVE, with reports of discriminative behaviors in various scenarios (Lippi-Green, 1997; Purnell et al., 1999; Massey and Lundy, 2001; Grogger, 2011; Rickford and King, 2016). We hire AAVE speakers to manually rewrite instances from seven popular SE reasoning benchmarks into AAVE (Section 2.1). Our approach has unique advantages compared with prior works which either (i) rely on predefined lexical or morphosyntactic transformation rules (Ziems et al., 2022, 2023), which may overlook subtle contextual nuances, or (ii) use LLM as translators (Gupta et al., 2024), which may have the very biases that our research wants to unveil (Fleisig et al., 2024; Smith et al., 2024).
Category Algorithm (25.7%) Logic (29.8%) Math (24.7%) Intergrated (19.7%) Total
@@
Table 1: ReDial contains $1,216$ fully-annotated parallel prompts for four categories, drawn from seven data sources. Each category contribution to the total amount is reported in percentage points in brackets.
We introduce ReDial (Reasoning with Dialect Queries), the first high-quality, end-to-end human-annotated SE-AAVE parallel dataset for reasoning tasks (Section 2). ReDial contains over 1.2K SE-AAVE prompt pairs covering four canonical reasoning categories: algorithm, math, logic, and integrated reasoning (tasks that require composing multiple reasoning skills). By anchoring these queries to known correct answers and employing human-based rewriting, ReDial furnishes an objective measure of dialect fairness and robustness. It also avoids the pitfalls of LLM-based evaluations, which can be inherently biased (Zheng et al., 2023; Chen et al., 2024; Shi et al., 2024).
Using ReDial, we benchmark widely used LLMs, including GPT-o1, GPT-4o, Claude-3.5-Sonnet, LLaMA-3.1-70B-Instruct, and others (Section 3). We find that almost all models experience statistically significant performance drops on AAVE prompts, despite their semantic equivalence to their SE counterparts. On average, we observe a relative performance reduction of more than 10%. This discrepancy persists even with advanced prompting techniques like Chain-of-Thought (CoT) prompting (Kojima et al., 2022; Wei et al., 2022), indicating that current LLMs are both brittle and unfair with dialectal inputs.
To understand these gaps, we further analyze potential causes. Our analysis reveals that the brittleness of LLMs with AAVE prompts arises from a combination of dialect-specific morphosyntactic features and nuanced conversational norms. Experiments with synthetic perturbations and AAVE-specific feature injections show that while these factors contribute to performance degradation, they fail to replicate the severity observed with human-annotated data. This highlights the limitations of rule-based transformations and the critical need for high-quality, context-rich datasets like ReDial to evaluate LLM fairness and robustness effectively.
In summary, in this paper:
  1. 1.
    We introduce ReDial, the first high-quality, end-to-end human-annotated AAVE-SE parallel reasoning benchmark spanning four foundational reasoning tasks.
  2. 2.
    We show that leading LLMs exhibit significant unfairness and brittleness on AAVE prompts compared to their SE counterparts.
  3. 3.
    We identify that the brittleness of LLMs with AAVE prompts stems from a combination of dialect-specific morphosyntactic features and nuanced conversational norms, which cannot be captured by synthetic transformations.
2 Dataset
ReDial (Reasoning with Dialect Queries) is a benchmark of more than 1.2K parallel Standard English – African American Vernacular English (SE-AAVE) query pairs. Table 1 provides an overview of the distribution, and Figure 1 along with Appendix A.2 present illustrative examples.
Following Zhu et al. (2023a), ReDial includes four canonical reasoning tasks—algorithm, logic, math. We additionally consider integrated reasoning as a compositional task requiring multiple skills. The task formulation of ReDial is linguistically diverse, addresses cornerstone problems in human reasoning, and is of particular interest as it is challenging for LLMs.
We first describe the data sources and sampling strategies (Section 2.1), and then detail the AAVE rewriting and validation processes that ensure high data quality (Section 2.2).
@@ 2.1 Data Sourcing
We construct a highly curated dataset by drawing upon seven established benchmarks covering different aspects of reasoning. For each source, we provide key references, task descriptions, and sample sizes. Additional examples can be found in Appendix A.1.
@@
Figure 2: Annotation and cross-validation of ReDial instances. We first sample instances from datasets of four canonical reasoning tasks to compose the source data, then we hire AAVE speakers to rewrite the instances in their dialect. To ensure the quality, we conduct naturalness check by AAVE speakers and correctness check by non-AAVE speakers and LLMs. We reannotate instances that do not pass the quality checks and iterate the process until the data meet our criteria. Finally, we combine the source data and AAVE rewriting to obtain ReDial.
Algorithm HumanEval (Chen et al., 2021) consists of 164 human-written code completion instances. We convert and include all these code completion headings into instruction-following natural language queries following the paradigm of InstructHumanEval. [^2] https://huggingface.co/datasets/codeparrot/instructhumaneval
Algorithm MBPP Austin et al. (2021) includes $1,000$ code generation queries. We randomly sample $150$ instances from its sanitized test set (Liu et al., 2023).
Math GSM8K (Cobbe et al., 2021) is a graduate-level math word problem dataset containing $8,790$ instances. We randomly sample $150$ instances from its test set.
Math SVAMP (Patel et al., 2021) is a collection of $1,000$ elementary-school math problems. We randomly sample $150$ instances from its test set.
Logic LogicBench (Parmar et al., 2024) comprises various logic questions in both binary classification and multiple-choice formats. We sample 100 binary and 100 multiple-choice questions, collecting $200$ samples in total.
Logic Folio (original+perturbed) (Han et al., 2022; Wu et al., 2023) Original Folio is a manually curated logic benchmark. We select $81$ instances along with their manually perturbed versions from Wu et al. (2023), yielding $162$ instances.
Intergrated AsyncHow (Lin et al., 2024) is a planning reasoning benchmark. LLMs must interpret natural language descriptions (i.e., logic), find different possible paths in the graph (i.e., algorithm), and then calculate and compare the time cost for these paths (i.e., math) to reach the correct answer. We use stratified sampling based on the dataset’s complexity metrics and include $240$ instances.
2.2 Annotation and Quality Assurance
After data sourcing, we hire AAVE speakers to rewrite each instance in AAVE. We schematize our annotation and validation in Figure 2 and describe them below.
Annotation.
We hire and ask AAVE speakers to rewrite each SE query so that it sounds natural to AAVE speakers while retaining all critical information (e.g., numerical values, logical conditions, and technical details). For algorithm-related tasks involving code, we hire annotators with a background in computer science to keep the logic and semantics of the code tasks. [^3] Details on annotator compensation, qualifications, and guidelines are presented in Appendix A.3.
Validation.
@@ 2.2 Annotation and Quality Assurance
We then perform a careful quality check to ensure both naturalness and correctness. First, we ask annotators to cross-check and edit each others’ annotations to make sure that the annotations are natural to AAVE speakers. Second, to ensure correctness, we first have non-AAVE speakers manually check the essential information, then conduct a sanity check with GPT-4o for the correctness of rewriting (details in Appendix A.4). We manually check data that GPT-4o flags as invalid to see if all essential information is preserved. No instance is rejected solely based on the LLM’s judgment. We return invalid instances to AAVE speakers for correction and iterate the process until all data pass the check.
After these efforts, we obtain ReDial: a high-quality, end-to-end human-annotated SE–AAVE parallel dataset comprising over 1.2K instances spanning four canonical reasoning tasks. In the rest of this paper, we refer to the SE portion of the dataset as SE ReDial and the AAVE portion as AAVE ReDial.
3 Experiments